OpenStudy (anonymous):

Solve the equation 9 = (square root of 27)^(4x + 6). Using complete sentences, explain the strategy used in solving this equation.

mathslover (mathslover):

Can you first write in the form of equation like using numeral values..?

OpenStudy (anonymous):

\[9=\sqrt{27}^{4x+6}\]

mathslover (mathslover):

Hmm okay, I will not be able to just write the whole strategy for you but I will explain you

OpenStudy (anonymous):

okay thanks

mathslover (mathslover):

Our first and foremost step will be to make each term in the powers of 3 so that we can equate the powers later on.

mathslover (mathslover):

Can you tell me, what is 9 in the powers of 3?

OpenStudy (anonymous):

is it 729?

mathslover (mathslover):

No, I meant that : 9 = 3^x What is x?

mathslover (mathslover):

9 in the powers of 3 means 9 in terms of 3 to the power of something if it was 9 to the powers of 3 then it would have been 729 but that is not the case for now.

OpenStudy (anonymous):

2

mathslover (mathslover):

Okay good, so : \(\color{blue}{\bf{9 = 3^2}}\)

mathslover (mathslover):

So, I will write the equation as : \(\large \color{blue}{\bf{3^2}} = \sf{\sqrt{27}^{4x+6}}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so then do you take the square root of 27?

mathslover (mathslover):

So, now we have to represent the Right Hand Side in the TERMS OF POWERS OF 3 : \(\large \sqrt{27}^{4x+6} = 3^y \) So, to calculate what is y, we have to do a lot of hardwork. Ready captain? :P

OpenStudy (anonymous):

yea I'm ready

mathslover (mathslover):

Good! One medal for that to you!

mathslover (mathslover):

Okay now, let us get back to work. \(\sqrt{27} \) , it can be written as : \(\sqrt{3 \times 3 \times 3 }\)

mathslover (mathslover):

So, it becomes : \(\sqrt{27} = \boxed{3\sqrt{3}}\) Right?

OpenStudy (anonymous):

right

mathslover (mathslover):

Okay so, we can right : \(\sqrt{3} \) as \(3^{\cfrac{1}{2}}\) Right?

mathslover (mathslover):

So, it becomes : \(3\sqrt{3} = 3 \times 3^\frac{1}{2} \) Can you simplify it?

OpenStudy (anonymous):

10 1/2

mathslover (mathslover):

Oh no.. See it is a theorem : \(\large a^n \times a^m = a^{m+n} \)

OpenStudy (anonymous):

im confused

mathslover (mathslover):

Okay, I will help you to get out from confusion.

mathslover (mathslover):

See, I have : \(\sqrt{27}\) as \(3\sqrt{3}\) = \(3^1 \times 3^\frac{1}{2}\) Now, as 3 is common I can write it as : \(3^{1 + \frac{1}{2}}\) What I did is just added up the powers as the base was same..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it would be \[3^{5/4}\]

mathslover (mathslover):

So, it is : \(\large{ 3^{1+ \cfrac{1}{2}}}\)

OpenStudy (anonymous):

typo sorry i mean 5/2

mathslover (mathslover):

No Shay. it is neither 5/2 nor 5/4 1 + 1/2 = (2+1)/2 = 3/2

mathslover (mathslover):

So, it will be : \(\large{3^\cfrac{3}{2}}\) Did you get it now?

OpenStudy (anonymous):

yea i did the adding wrong

mathslover (mathslover):

Okay, no problem, it happens!

mathslover (mathslover):

So, what we have now is : \(\sqrt{27} = 3^{\cfrac{3}{2}}\)

OpenStudy (anonymous):

okay

mathslover (mathslover):

So, plugin this value in the given equation : 9 = \(\sqrt{27} ^{4x+6}\) Can you tell me what we will have the equation as now?

mathslover (mathslover):

Just plugin the values of 9 and \(\sqrt{27}\) that we found above in the equation/.

OpenStudy (anonymous):

\[3^{2}=3\frac{ 3 }{ 2 }^{4x+6}\]

mathslover (mathslover):

Okay well it will be : \(3^2 = 3^{\cfrac{3}{2}(4x+6)}\)

mathslover (mathslover):

So, another theorem in use comes here : Whenever the base is same , just equate the powers like : \(a^m = a^n\) then m = n

mathslover (mathslover):

So here 3 is same in both the terms. So, we will just equate the powers. \(2 = \cfrac{3}{2} (4x+6)\) Is it okay?

OpenStudy (anonymous):

no i don't think so

mathslover (mathslover):

Hmm, where are you confused?

OpenStudy (anonymous):

i don't get how the three got eliminated

mathslover (mathslover):

Well, it is a theorem and its' proof is a little bit higher level.

mathslover (mathslover):

The theorem says that : if \(\LARGE{\boxed {\color{Red}{a}^{\color{blue}{m}} = \color{red}{a}^{\color{orange}{n}}}}\) Then: \(\LARGE{\boxed {\color{blue}{m} = \color{orange}{n}}}\) You can say, we didn't eliminate a from there , in fact we just took the powers from there and we are solving that.

mathslover (mathslover):

(If you have studied logarithm , then this is the proof: ) \(a^m = a^n\) Taking log both sides : \(\large{ \sf{ \log {a^m} = \log {a^n} \\ m \log a = n \log a }\\ \textbf{ log a cancels out both sides and we get :} \\ m = n }\)

mathslover (mathslover):

I know, you have not studies logarithm yet, so it will be very difficult to you to understand that but for now, just understand that when base is same ( in this case , it is 3 ) then the powers are also equal.

mathslover (mathslover):

Are you still confused?

OpenStudy (anonymous):

okay. no not confused

mathslover (mathslover):

Great . So, you have now : \(2 = \cfrac{3}{2} (4x+6)\) Right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

So, isn't it easy to solve for x now? Can you do this Captain?

OpenStudy (anonymous):

would you foil?

mathslover (mathslover):

Yes you can use it here or just to make it easier with the calculation : \(2 = \cfrac{3}{2} (4x+6)\) Take 2 common from 4x + 6 \(2 = \cfrac{3}{2} (2(2x+3))\)

mathslover (mathslover):

So, : \(\sf{ 2 = \cfrac{3}{\cancel{2}} \times \cancel{2} (2x+3)}\)

mathslover (mathslover):

So, we get finally : 2 = 3(2x+3) It will be easier now to find x. Can you do that?

OpenStudy (anonymous):

yes

mathslover (mathslover):

Okay, so, tell me what you get the value for x

OpenStudy (anonymous):

x=1.16

OpenStudy (anonymous):

*x=-1.16

mathslover (mathslover):

yes it is correct.

mathslover (mathslover):

Or you can say : \(x = -\cfrac{7}{6}\)

OpenStudy (anonymous):

Thank you!

mathslover (mathslover):

Thanks for the medal. And Good Luck for your journey on OpenStudy!

OpenStudy (anonymous):

Thanks!